OpenStudy (lgbasallote):

\[\mathcal L \lbrace t^{5/2} \rbrace\]

OpenStudy (lgbasallote):

i got \[\frac{5/2 \Gamma (5/2)}{s^{5/2+1}}\] \[\implies \frac{(5/2)(3\sqrt \pi / 4)}{s^{7/2}}\] \[\implies \frac{15\sqrt \pi/8}{s^3 \sqrt s}\] \[\implies \frac{15}{8s^3} (\frac{\pi}{s})^{1/2}\] is this right?

OpenStudy (experimentx):